A man invests a total of $9,493in two savings accounts. One account yields 9% simple interest and the other 10% simple interest. He earned a total of $930.80interest for the year. How much was invested in the 9% account?

Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 10.00% per annum ------------ Amount invested = y
9493
Interest----- 930.80

Part I 9.00% per annum ---x
Part II 10.00% per annum ---y
Total investment
x + 1 y= 9493 -------------1
Interest on both investments
9.00% x + 10.00% y= 930.8
Multiply by 100
9 x + 10 y= 93080.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -85437.00
Add this to (2)
0 x 1 y= 7643
divide by 1
y = 7643
Part I 9.00% $ 1850
Part II 10.00% $ 7643

CHECK
1850 --------- 9.00% ------- 166.50
7643 ------------- 10.00% ------- 764.30
Total -------------------- 930.80

m.ananth@hotmail.ca